OpenStudy (anonymous):

What would be the standard form of this equation? 3/4y-2=1/2x

OpenStudy (anonymous):

3/4y-2=1/2x 3y-8=2x -2x+3y=8

OpenStudy (anonymous):

ok?

OpenStudy (lgbasallote):

i think it's written as \[\frac{3}{4y - 2} = \frac{1}{2x}\] but if it is written \frac{3}{4} y - 2 = \frac{1}{2} x\] then @cyclevleoce12345 is right

OpenStudy (lgbasallote):

but if it is written \[\frac{3}{4} y - 2 = \frac{1}{2} x\] *